OpenStudy (anonymous):

Conditional Probability: A city council consists of 6 Democrats, two of whom are women, and 6 republicans, 4 of whom are men. A member is chosen at random. If the member chosen is a man, what is the probability that he is a democrat?

jimthompson5910 (jim_thompson5910):

6 democrats: 2 women, 4 men 6 republicans: 4 men, 2 women

jimthompson5910 (jim_thompson5910):

how many men are there total

OpenStudy (anonymous):

i think just 4

jimthompson5910 (jim_thompson5910):

there are 4 men (democrats) + 4 men (republicans) = 8 men total

jimthompson5910 (jim_thompson5910):

so the probability of picking a democrat, given the choice is a male, is 4/8 = 1/2

OpenStudy (anonymous):

oh thank you

jimthompson5910 (jim_thompson5910):

yw